# Free Electron Paths from Dirac’s Wave Equation Elucidating Zitterbewegung and Spin

### v2

@@ 3.1 Neoclassical and Dirac’s Energy Equations
Beck (2023) derived an energy equation based on Eq. (3.8) that extends the usual “on-shell” relativistic energy-momentum equation by an additional term arising from the electron’s magnetic and electric dipole energies:
@@
using his derived equation $\pi^{\mu}z_{\mu}=0$. The orthogonal 3-vectors $\mathbf{s}=\mathbf{z}\times m\mathbf{u}$ and $\mathbf{d}=m(u^{0}\mathbf{z}-z^{0}\mathbf{u})$ are defined in the usual way by the space and time components of the spin tensor $S^{\mu\nu}$ defined by Eq. (3.6). The physics behind these two vectors is more transparent when they are expressed relative to the rest-frame $X_{c}$ as $\mathbf{s}=\mathbf{r}\times m\dot{\mathbf{r}}$ and $\mathbf{d}=mc\mathbf{r}$. Notice that the sum of the dipole energies is $\Phi(\tau)=\frac{1}{2}P^{\mu\nu}(\tau)F_{\mu\nu}(x(\tau))$, where $P^{\mu\nu}=-\frac{q}{m}S^{\mu\nu}$ is the dipole moment tensor, explicitly showing the dipole coupling to the em-field.
@@
where the kinematic momentum operator $\hat{\pi}_{\mu}=i\hbar\partial_{\mu}-qA_{\mu}$ with $A_{\mu}$ the covariant components of the 4-vector em-potential, and $\hat{\Phi}$ is the dipole operator corresponding to the interaction energy from the electric and magnetic dipoles set up by the spin motion of the electron
@@ 3.2 Gordon Decomposition for a Free Electron
As shown in Beck (2023) in his Eq. (2.15), $z^{\mu}\pi_{\mu}=0$, implying by differentiation that $\dot{z}^{\mu}\pi_{\mu}=0$ since $\dot{\pi}_{\mu}=0$ for a free electron. It then follows from the definition of $\Sigma^{\mu\nu}$ that $\Sigma^{\mu\nu}\pi_{\nu}=0$. It is also shown in Beck’s Eq. (3.11)(iii) that $S^{\mu\nu}\pi_{\nu}=-(mc)^{2}z^{\mu}$, which implies by differentiation that $\dot{S}^{\mu\nu}\pi_{\nu}=-(mc)^{2}\dot{z}^{\mu}$. Substituting these results into Eq. (3.16)
@@
The polarization vector is $\frac{1}{mc}\mathbf{d}=\dot{t}\mathbf{z}-t_{z}\mathbf{u}$, which is just the zitter radius vector $\mathbf{r}$ in the rest-frame $X_{c}$ where $\mathbf{y}=0$ and so it is the electron’s spatial position $\mathbf{x}$ relative to $X_{c}$. The spatial part $q\widetilde{\mathbf{w}}(x)$ is the sum of polarization and magnetization currents (e.g. Baym (1981)). Therefore, these currents are a consequence of the zitter motion of the electron. It is a curious result that the magnetization current $\frac{q}{m}\nabla\times\mathbf{s}$ is zero in the rest-frame for a free electron since the spin vector $\mathbf{s}=\bar{\psi}(x)\hat{\mathbf{s}}\psi(x)$ is constant there. However, the circular zitter motion in the rest-frame is still present, of course, as are the magnetic and electric dipole moment 3-vectors, $\mu$ $\frac{q}{m}\mathbf{s}$ and $\epsilon$ $\frac{q}{mc}\mathbf{d}$ ($q\mathbf{r}$), respectively. The zitter motion then corresponds entirely to the polarization current, which is from Eq. (3.18), $q\mathbf{w}^{\mu}=\frac{q}{mc}\frac{\partial}{\partial t}\mathbf{d}=q\dot{\mathbf{r}}^{\mu}$ since $t=\tau$ in the rest-frame. The spin axis has to precess for the magnetization current to be non-zero, as it does for a free electron when viewed by an observer moving relative to the electron’s rest-frame.
@@ 4 Concluding Remarks
In this work, we show for the first time that the classical particle model of the Dirac electron due to to Barut and Zanghi (1984) that uses “classical” spinors and classical dynamic variables of a proper time, is equivalent to Dirac’s equation in the case of a free electron, that is, every Barut-Zanghi spinor solution is a Dirac wave function solution, and conversely. It has already been shown in Beck (2023) that there is an equivalence between the equations of motion in classical dynamic variables that are derived by applying Dirac’s velocity and spin operators to the Barut-Zanghi spinor, and the equations of motion of Beck’s neoclassical relativistic mechanics model of the electron. In addition, it is noted in this work that the classical particle models of Hestenes (2010) and Barut and Zanghi (1984) that both use spinors to express the dynamics of an electron are equivalent, and that the classical spin models of Beck (2023), Rivas (2003) and Salesi (2002) are also equivalent. Therefore, all these classical models describe the same dynamics for the spinning electron, and, for the free electron, they all agree with the dynamics expressed by Dirac’s wave equation.
@@
Furthermore, since Dirac’s equation for a free electron implies the equations of motion of the neoclassical model presented by Beck (2023), his previous conclusions also follow from Dirac’s equation, such as
@@
One obvious challenge to extending this work for an electron in an em-field is that, unlike the free electron case, there is no analytical solution $\psi(x)$ of Dirac’s equation, in general, although there is for a uniform magnetic field, suggesting a fruitful direction for future research. When an analytical solution is available, Dirac’s velocity and spin operators can be applied to the wave function to evaluate the proper velocity $\widetilde{u}^{\mu}(x)$ for the electron at space-time point $x$, along with the spin tensor $\widetilde{S}^{\mu\nu}(x)$. The Gordon decomposition in Eq. (2.10) should then give the inertia center velocity $\widetilde{v}^{\mu}(x)=\widetilde{\dot{y}}^{\mu}(x)$ and the zitter (spin) velocity $\widetilde{w}^{\mu}(x)=\widetilde{\dot{z}}^{\mu}(x)$. Another challenge is that the relation between the Dirac wave function phase $\theta(x)$ and the proper time $\tau(x)$ in Eq. (2.12) must be expressed in incremental form and integrated along the electron’s worldline to get $\tau(x)$.
It also remains to be determined if the approach can be extended to $N$ interacting electrons. One challenge in this case is that proper time is specific to an electron and so there will be $N$ proper times, each of which must be related to an inertial observer’s time. Each electron has two worldlines: one for its position in proper time (the charge location) and one for its inertia (or spin) center. In the classical Dirac particle models, these two worldlines are described by two coupled second-order ordinary differential equations and it remains to be seen if that is also true for Dirac’s equation.

### v3

@@ 3.1 Neoclassical and Dirac’s Energy Equations
Beck derived an energy equation based on Eq. (3.8) that extends the usual “on-shell” relativistic energy-momentum equation by an additional term arising from the electron’s magnetic and electric dipole energies:
@@
using his derived equation $\pi^{\mu}z_{\mu}=0$ (Beck (2023)). The orthogonal 3-vectors $\mathbf{s}=\mathbf{z}\times m\mathbf{u}$ and $\mathbf{d}=m(u^{0}\mathbf{z}-z^{0}\mathbf{u})$ are defined in the usual way by the space and time components of the spin tensor $S^{\mu\nu}$ defined by Eq. (3.6). The physics behind these two vectors is more transparent when they are expressed relative to the rest-frame $X_{c}$ as $\mathbf{s}=\mathbf{r}\times m\dot{\mathbf{r}}$ and $\mathbf{d}=mc\mathbf{r}$. Notice that the sum of the dipole energies is $\Phi(\tau)=\frac{1}{2}P^{\mu\nu}(\tau)F_{\mu\nu}(x(\tau))$, where $P^{\mu\nu}=-\frac{q}{m}S^{\mu\nu}$ is the dipole moment tensor, explicitly showing the dipole coupling to the em-field through the work done by the field on the dipole current.
@@
where the kinematic momentum operator is $\hat{\pi}_{\mu}=i\hbar\partial_{\mu}-qA_{\mu}$ with $A_{\mu}$ the covariant components of the 4-vector em-potential, and $\hat{\Phi}$ is the dipole operator corresponding to the interaction energy from the electric and magnetic dipoles set up by the spin motion of the electron
@@ 3.2 Gordon Decomposition for a Free Electron
As shown in Beck (2023) in Eq. (2.15), $z^{\mu}\pi_{\mu}=0$, implying by differentiation that $\dot{z}^{\mu}\pi_{\mu}=0$ since $\dot{\pi}_{\mu}=0$ for a free electron. It then follows from the definition of $\Sigma^{\mu\nu}$ that $\Sigma^{\mu\nu}\pi_{\nu}=0$. It is also shown in Beck’s Eq. (3.11)(iii) that $S^{\mu\nu}\pi_{\nu}=-(mc)^{2}z^{\mu}$, which implies by differentiation that $\dot{S}^{\mu\nu}\pi_{\nu}=-(mc)^{2}\dot{z}^{\mu}$. Substituting these results into Eq. (3.16)
@@
The polarization vector is $\frac{1}{mc}\mathbf{d}=\dot{t}\mathbf{z}-t_{z}\mathbf{u}$, which is just the zitter radius vector $\mathbf{r}$ in the rest-frame $X_{c}$ where $\mathbf{y}=0$ and so it is the electron’s spatial position $\mathbf{x}$ relative to $X_{c}$. The spatial part $q\widetilde{\mathbf{w}}(x)$ is the sum of polarization and magnetization currents (e.g. Baym (1981)). Therefore, these currents are a consequence of the zitter motion of the electron. It is a curious result that the magnetization current $\frac{q}{m}\nabla\times\mathbf{s}$ is zero in the rest-frame for a free electron since the spin vector $\mathbf{s}=\bar{\psi}(x)\hat{\mathbf{s}}\psi(x)$ is constant there. However, the circular zitter motion in the rest-frame is still present, of course, as are the magnetic and electric dipole moment 3-vectors, $\mu$ $\frac{q}{m}\mathbf{s}$ and $\epsilon$ $\frac{q}{mc}\mathbf{d}$ ($q\mathbf{r}$), respectively. The zitter motion then corresponds entirely to the polarization current, which is from Eq. (3.18), $q\mathbf{w}=\frac{q}{mc}\frac{\partial}{\partial t}\mathbf{d}=q\dot{\mathbf{r}}$ since $t=\tau$ in the rest-frame. The spin axis has to precess for the magnetization current to be non-zero, as it does for a free electron when viewed by an observer moving relative to the electron’s rest-frame.
@@ 4 Concluding Remarks
In this work, we show for the first time that the classical particle model of the Dirac electron in Barut and Zanghi (1984) that uses “classical” spinors and classical dynamic variables of a proper time, is equivalent to Dirac’s equation in the case of a free electron, that is, every Barut-Zanghi spinor solution is a Dirac wave function solution, and conversely. It has already been shown in Beck (2023) that there is an equivalence between the equations of motion in classical dynamic variables that are derived by applying Dirac’s velocity and spin operators to the Barut-Zanghi spinor, and the equations of motion of Beck’s neoclassical relativistic mechanics model of the electron. In addition, it is noted in this work that the classical particle models in Hestenes (2010) and Barut and Zanghi (1984) that both use spinors to express the dynamics of an electron are equivalent, and that the classical spin models in Beck (2023), Rivas (2003) and Salesi (2002) are also equivalent. Therefore, all these classical models describe the same dynamics for the spinning electron, and, for the free electron, they all agree with the dynamics expressed by Dirac’s wave equation.
@@
Furthermore, since Dirac’s equation for a free electron implies the equations of motion of the neoclassical model presented in Beck (2023), his previous conclusions also follow from Dirac’s equation, such as
@@
One obvious challenge to extending this work for an electron in an em-field is that, unlike the free electron case, there is no analytical solution $\psi(x)$ of Dirac’s equation, in general, although there is for a uniform magnetic field, suggesting a fruitful direction for future research. When an analytical solution is available, Dirac’s velocity and spin operators can be applied to the wave function to evaluate the proper velocity $\widetilde{u}^{\mu}(x)$ for the electron at space-time point $x$, along with the spin tensor $\widetilde{S}^{\mu\nu}(x)$. The Gordon decomposition in Eq. (2.10) should then give the inertia center velocity $\widetilde{v}^{\mu}(x)=\widetilde{\dot{y}}^{\mu}(x)$ and the zitter (spin) velocity $\widetilde{w}^{\mu}(x)=\widetilde{\dot{z}}^{\mu}(x)$.
Another challenge in this case is that the relation between the Dirac wave function phase $\theta(x)$ and the proper time $\tau(x)$ in Eq. (2.12) must be expressed in incremental form and integrated along the electron’s worldline to get $\tau(x)$, so they must be established simultaneously. In Beck’s classical Dirac particle model covered in this paper, this worldline, which gives the charge position in proper time, is governed by a fourth-order ordinary differential equation that can be expressed as two coupled second-order ordinary differential equations that also includes the worldline for the electron’s inertia (or spin) center. It remains to be seen if that is also true for Dirac’s equation when an em-field is present.
It also remains to be determined if the approach can be extended to $N$ interacting electrons. One challenge in this case is that proper time is specific to an electron and so there will be $N$ proper times, each of which must be related to an inertial observer’s time. Of course, for this case, Dirac’s single-electron equation is usually replaced by a quantum field theory interpretation of it.